OpenStudy (anonymous):

Find a ⋅ b. a = 4i - 4j, b = 4i + 5j

jimthompson5910 (jim_thompson5910):

if x = ai+bj y = ci + dj then x dot y = a*c + b*d

jimthompson5910 (jim_thompson5910):

Basically, to find the dot product of two vectors, you step 1) multiply the corresponding components step 2) add up the products found in step 1

OpenStudy (anonymous):

So it would come out to be -36?

jimthompson5910 (jim_thompson5910):

That is incorrect.

OpenStudy (anonymous):

ah -4. Did my math wrong.

jimthompson5910 (jim_thompson5910):

much better